how to graph fractions like 5/6,4/3 on a number line
Answer by richwmiller(17219)

either put more tick marks between whole numbers or guess.
4/3=8/6 =1 2/6
5/6=5/6
or convert to decimals
You have just discovered one of the weaknesses of using graphs to solve equations.
They are hard to read and write if the data is not a whole number.